# NRSeg: Noise-Resilient Learning for BEV Semantic Segmentation via Driving World Models

### v1

@@ Abstract
Index Terms:
Bird’s-eye-view understanding, driving world model, unsupervised domain adaptation, evidential deep learning.
@@ I Introduction
To improve BEV understanding in real-world scenes, Unsupervised Domain Adaptation (UDA) and Semi-Supervised Learning (SSL) have become prevalent approaches for deploying perception tasks in unlabeled environments. In BEV semantic segmentation, numerous studies have explored these two methodologies. Specifically, DualCross [8] employed adversarial learning to achieve UDA, while PCT [3] introduced perspective pseudo-label supervision for both UDA and SSL. SemiBEV [9], designed for single-view tasks, enhanced SSL efficiency through a spatially-rotated data augmentation strategy. Although these methods have improved the accuracy of UDA and SSL in BEV perception, they fundamentally rely on existing annotated data, thereby achieving only limited improvements in learning capability. Furthermore, since the BEV semantic segmentation task is highly data-hungry, with its performance heavily dependent on the scale of available labeled data, enhancing labeled training data can significantly improve model performance. However, the annotation process of BEV labels remains highly labor-intensive. The rapid advancement of generative models [10, 11, 12] has attracted our attention, as these models can synthesize highly diverse and photorealistic images from labels. Thus, recent advances in driving world models [13, 2, 1], which generate multi-view street images from BEV semantic labels, vehicle bounding boxes, and text prompts, have raised an important yet underexplored question: Does incorporating diversely distributed synthetic images with BEV labels improve BEV semantic perception performance?
@@
To enhance the noise robustness of BEV models, this work proposes a novel paradigm, NRSeg, a noise-resilient learning framework for BEV semantic segmentation, as shown in Fig. LABEL:Fig.intro_new(a). The proposed method simultaneously explores optimized guidance for synthetic data of varying quality and robust learning of the model itself against disturbances. A Perspective-Geometry Consistency Metric (PGCM) is designed to quantitatively assess the contribution of synthetic data to model learning. Under perspective view, the contribution scores of synthetic data are quantified by evaluating the similarity between road masks in synthetic images and their BEV back-projected counterparts. To alleviate the effect of generation noise in synthetic data, this score adaptively encourages the model to focus on non-labeled regions of synthetic data by guiding the loss optimization direction. Furthermore, the model’s robustness against disturbances is enhanced through dual-distribution constraints on predictions in the Bi-Distribution Parallel Prediction (BiDPP) module. While the multinomial distribution directly predicts semantic probabilities of a single pixel, the Dirichlet distribution is employed to model the uncertainty relationships based on Evidential Deep Learning (EDL) [14, 15]. The premise of EDL requires mutually exclusive semantic categories, which cannot be satisfied in BEV tasks. To effectively adapt EDL for BEV semantic segmentation tasks, the Hierarchical Local Semantic Exclusion (HLSE) module is designed, which partitions locally mutually exclusive semantic categories into unit clusters and achieves uncertainty modeling through hierarchical fusion of their independent learning processes.
@@ II-A BEV Semantic Segmentation
BEV semantic mapping performs semantic segmentation of static objects in road environments from a Bird’s Eye View (BEV). The fundamental challenge of BEV semantic segmentation lies in inferring 3D information from 2D inputs. Depending on the approach to view transformation, this research currently encompasses three distinct methodological categories. One dominant approach, exemplified by the LSS framework [6, 17], achieved spatial transformation through 2D pixel depth prediction. The alternative paradigm adaptively learned spatial feature representations by projecting 3D sampling points onto 2D image planes [18, 19]. Furthermore, some approaches [20, 7] leveraged the prior knowledge that static road objects predominantly lie on the ground plane, utilizing this prior information to facilitate view transformation through Inverse Perspective Mapping (IPM). While these view transformation methods demonstrate efficient adaptation to diverse road environments, their performance significantly degrades in complex, congested scenarios due to the sparsity of 2D observations. To address this, StreamMapNet [21] introduced a streaming temporal strategy [22, 23] that compensated for single-frame sparsity by aggregating long-term historical information, improving mapping accuracy.
@@
Although these methods demonstrate competitive performance across domains, these methods have not yet explored directly enhancing the learning capability of the model by increasing the diversity of data distributions within the source domain, where significant accuracy improvements remain achievable. Benefiting from the rapid advancement of generative models [25, 26], it is now feasible to utilize simple yet effective information to synthesize diverse data. We aim to explore the potential of synthetic data from driving world models for robustifying BEV semantic segmentation.
@@ III-A Framework of the NRSeg Model
$Loss_{pv}=\text{SegLoss}(P_{v},L_{v}).\vspace{-0.3em}$ (1)
@@
$\vspace{-0.3em}F_{b}^{t^{\prime}}=\text{TFu}(\text{Ailgh}(F_{b}^{t-1}),F_{b}^{t}).$ (2)
@@
Moreover, $N$ is designed to contain a shared BEV encoder and two segmentation detectors: a multinomial distribution $D_{m}^{k}$ and a Dirichlet distribution $D_{d}^{k}$ for predictions.
$\vspace{-0.5em}D_{m}^{k},D_{d}^{k}=N(F_{b}^{t^{\prime}}),$ (5)
where $k$ is the number of semantic classes. The semantic probability predictions can be obtained by $P_{m}^{k}=\text{Sigmoid}(D_{m}^{k})$. Typically, the BEV semantic segmentation models are supervised by semantic labels $L_{b}$, which guide model optimization by constraining the predicted semantic probabilities $P_{m}^{k}$. Following the work of [9], the DICE loss is adopted:
@@
In contemporary BEV perception datasets, perspective images and semantic labels are precisely aligned. Yet, due to generation errors in world models, synthetic data and semantic labels cannot maintain perfect correspondence. As shown in Fig. 1, to mitigate the misleading effects of such misalignment, this paper proposes Perspective-Geometry Consistency Metric (PGCM) and Bi-Distribution Parallel Learning (BiDPP) to strengthen noise-resilient learning and model robustness in BEV semantic segmentation.
@@ III-B Perspective-Geometry Consistency Metric
First, the BEV semantic label and 3D object boxes are projected onto the perspective view space, maintaining consistency with the process described in works [44, 24, 1]. Assuming the BEV label lies on the ground plane, the height is $h=0$. With this prior knowledge, each cell $(X,Y)$ of the BEV label $M_{b}$ can be projected into perspective view $(u,v)$:
$\vspace{-0.3em}[u,v,1]^{\text{T}}=\text{Proj}_{\text{b2p}}(T_{\text{ex}},T_{\text{in}},[X,Y,h]^{\text{T}}),$ (7)
where $T_{ex}$ and $T_{in}$ is the camera parameters. $Proj_{b2p}(\cdot)$ is the conversion from BEV coordinates to perspective coordinates. Then the projected road mask of perspective view $M_{o}$ can be obtained. The 3D object boxes are directly back-projected onto the perspective image through Eq. 7 to obtain the projected object mask $M_{c}$. The perspective view is a 2D representation, inherently suffering from occlusion issues due to viewpoint limitations. Since targets like vehicles exist above the ground plane, they occlude the road surface during perspective imaging. Consequently, the final reference mask should adhere to the following criteria:
$\vspace{-0.5em}M_{r}=M_{o}*\overline{M_{c}},$ (8)
where $\overline{M_{c}}$ is the negation of $M_{c}$. For perspective masks of synthetic images $M_{s}$, they can be produced using generalized semantic segmentation models such as Mask2Former [45] and SAN [46].
To achieve accurate correspondence between synthetic data and BEV labels, the projected and synthetic perspective road masks should inherently maintain geometric consistency. In other words, any misalignment between them indicates inherent discrepancies between the synthetic map and BEV semantic labels. Therefore, the alignment error between the reference perspective masks $M_{r}$ and the synthetic perspective masks $M_{s}$ is used to quantify the BEV accuracy of the generated images. Considering that Intersection over Union (IoU) is a well-established metric for mask alignment evaluation, we adopt it as the primary measure here. The score of each frame oriented from synthetic data can be calculated by:
$\vspace{-0.5em}\textbf{R}=\frac{1}{n}\sum_{i=0}^{n}\text{IoU}(M_{r}^{i},M_{s}^{i}),$ (9)
where $n$ is the number of multi-view images in each frame. Then, for misaligned synthetic data samples and labels, this consistency score is incorporated as a weighting factor to guide the loss optimization direction in deep learning model training.
We employ the DICE semantic segmentation loss [47] as the learning objective to enable robust learning with noisy samples. The DICE loss is defined as:
@@
where $H$ and $W$ are the shapes of labels. $P$ is the prediction probabilities and $L$ is the labels. Since the labels follow a binary $0-1$ distribution, Eq. 10 can be simplified as:
$\vspace{-0.5em}loss_{gt}=1-\frac{2*P_{+}}{P_{+}+P_{-}+N_{+}+\sigma},$ (11)
@@
$\vspace{-0.5em}\min_{p+,p_{-}}1-\frac{2*P_{+}}{P_{+}+P_{-}+N_{+}+\sigma}.$ (12)
Intuitively, this requires the numerator and denominator of the right-hand term in Eq. 12 to be equal. In other words, this objective enforces two conditions: (i) $P_{+}\approx N_{+}$, and (ii) $P_{-}\approx 0$.
@@
$\vspace{-0.5em}loss_{gt}=1-\frac{2*P_{+}}{\textbf{R}*(P_{+}+P_{-}+N_{+}+\sigma)}.$ (13)
@@
$\vspace{-0.5em}2*P_{+}\approx\textbf{R}*(P_{+}+P_{-}+N_{+}+\sigma).$ (14)
@@
$\vspace{-0.5em}P_{-}\approx\frac{(2-\textbf{R})}{\textbf{R}}{\times}P_{+}-N_{+}-\sigma.$ (15)
@@ III-C Bi-Distribution Parallel Learning
In semantic segmentation tasks, the efficient multinomial distribution is typically adopted to predict semantic probabilities, where loss functions are utilized as constraints to assign a semantic category to each pixel. This result fundamentally relies on the absolute reliability of the multinomial distribution assumption. In practice, however, the multinomial distribution exhibits inherent uncertainty when handling ambiguous categories, leading to potentially unreliable predictions. Specifically, the synthetic data referenced in this study inherently contains noise, which significantly increases the likelihood of unreliable predictions. Therefore, this section additionally introduces the Dirichlet distribution with Evidential Deep Learning (EDL) to further analyze uncertainty in semantic predictions. In summary, this section realizes a bi-distribution parallel learning module, ensuring the efficiency and robustness of learning.
Rooted in Dempster-Shafer Theory (DST) and Subjective Logic, EDL formalizes belief assignments through Dirichlet distributions, enabling principled uncertainty quantification [14]. Specifically To learn a parametric representation of belief mass from prediction distributions, the Dirichlet distribution is employed as the prior for parameterization as $[\alpha_{0},...,\alpha_{k}]$:
@@
where $e_{k}=\text{ReLU}(D_{d}^{k})$ is the evidence information, which is derived from the distribution $D_{d}^{k}$ learned by the network. The parameterized distribution represents the density of probabilistic predictions $P_{e}$:
@@
where $S_{k}$ is the $k$-dimensional unit simplex. $B(\alpha)$ is a k-dimensional polynomial beta equation. Theoretically, the expected probability for the class corresponds to the mean of the Dirichlet distribution:
@@
where $S=\sum_{k}\alpha_{k}$. However, this theory fundamentally requires the mutual exclusivity of semantic categories per pixel. Assuming two semantic classes are non-exclusive, $\alpha_{1}\in(0,+\infty)$ and $\alpha_{2}\in(0,+\infty)$, their Dirichlet distribution means are:
@@
2:  $\alpha_{k}=\text{ReLU}(D_{d}^{k})+1$
3:  $\alpha=[\alpha_{0},\dots,\alpha_{k}]$
@@
Specifically, semantic classes are first hierarchically grouped based on their inherent properties, forming multi-level local clusters:
@@
Here, $n$ is the number of semantic classes in this cluster $G_{m}$ ($n<k$). Note that $n$ is not fixed to ensure adaptability across diverse tasks. Then, parameterization is implemented using the Dirichlet distribution as the prior within each local cluster, $G_{m}=[\alpha^{0},...,\alpha^{c_{n}}]$. Eq. 19 and Eq. 20 are executed to generate both predictive semantic probabilities and uncertainty measures. To train this head, the type-II maximum likelihood version of the uncertainty-aware loss is adopted in the local clusters:
$Loss_{e}^{G_{m}}=\sum_{i=0}^{H_{b}{\times}W_{b}}\sum_{j=0}^{C_{n}}L_{j}^{i}\cdot log(S/\alpha_{j}^{i}).$ (24)
@@
Boston $\xrightarrow{}$ Singapore
Source Only $128{\times}352$ 40.2 6.3 7.0 1.9 0.7 9.7 10.9
DualCross [8] $128{\times}352$ 43.8 8.2 13.5 4.5 3.6 9.6 13.9
DomainADV* [3] $224{\times}480$ 40.0 8.3 11.7 4.5 2.2 11.6 13.1
PCT* [3] $224{\times}480$ 46.2 8.6 14.2 6.4 3.7 15.0 15.7
MT+PV $128{\times}352$ 49.3 9.8 12.9 4.9 4.5 13.7 15.8
Ours $128{\times}352$ 59.1 16.9 21.9 12.1 16.8 21.0 24.6
Singapore $\xrightarrow{}$ Boston
Source Only $128{\times}352$ 44.7 2.6 11.5 4.2 0.2 8.4 11.9
DualCross [8] $128{\times}352$ 45.7 2.2 13.6 6.4 5.2 9.9 13.8
DomainADV* [3] $224{\times}480$ 35.7 4.2 11.3 4.8 0.6 9.7 11.1
PCT* [3] $224{\times}480$ 47.0 8.0 19.3 6.3 0.7 13.7 15.8
MT+PV $128{\times}352$ 48.0 2.3 16.8 7.8 4.9 10.5 15.0
Ours $128{\times}352$ 61.4 21.7 31.8 14.1 24.1 19.8 28.8
$\xrightarrow{}$ Rain
Source Only $128{\times}352$ 67.1 29.5 35.8 23.4 24.6 25.1 34.2
DualCross [8] $128{\times}352$ 67.9 29.8 38.4 22.6 29.3 25.4 35.6
DomainADV* [3] $224{\times}480$ 72.0 39.8 42.0 33.7 38.9 33.6 43.3
PCT* [3] $224{\times}480$ 78.3 45.2 52.1 37.6 47.2 36.4 49.5
MT+PV $128{\times}352$ 70.5 34.1 39.9 27.1 29.2 27.8 38.1
Ours $128{\times}352$ 72.8 36.6 42.5 29.8 41.7 30.4 42.3
Day $\xrightarrow{}$ Night
Source Only $128{\times}352$ 32.8 2.2 4.3 4.4 0.0 9.2 8.8
DualCross [8] $128{\times}352$ 49.4 8.0 12.7 7.9 0.0 15.1 15.5
DomainADV* [3] $224{\times}480$ 37.1 16.4 10.7 5.7 0.0 11.2 15.1
PCT* [3] $224{\times}480$ 51.3 19.4 16.1 7.6 0.0 19.3 19.0
MT+PV $128{\times}352$ 57.1 9.6 13.8 6.2 0.0 18.9 17.6
Ours $128{\times}352$ 61.1 31.3 18.6 12.0 0.0 22.3 24.2

### v2

@@ I Introduction
To improve BEV understanding in real-world scenes, Unsupervised Domain Adaptation (UDA) and Semi-Supervised Learning (SSL) have become prevalent approaches for deploying perception tasks in unlabeled environments. In BEV semantic segmentation, numerous studies have explored these two methodologies. Specifically, DualCross [25] employed adversarial learning to achieve UDA, while PCT [14] introduced perspective pseudo-label supervision for both UDA and SSL. SemiBEV [53], designed for single-view tasks, enhanced SSL efficiency through a spatially-rotated data augmentation strategy. Although these methods have improved the accuracy of UDA and SSL in BEV perception, they fundamentally rely on existing annotated data, thereby achieving only limited improvements in learning capability. Furthermore, since the BEV semantic segmentation task is highly data-hungry, with its performance heavily dependent on the scale of available labeled data, enhancing labeled training data can significantly improve model performance [24]. However, the annotation process of BEV labels remains highly labor-intensive. The rapid advancement of generative models [46, 27, 33] has attracted our attention, as these models can synthesize highly diverse and photorealistic images from labels. Thus, recent advances in driving world models [45, 9, 50], which generate multi-view street images from BEV semantic labels, vehicle bounding boxes, and text prompts, have raised an important yet underexplored question: Does incorporating diversely distributed synthetic images with BEV labels improve BEV semantic perception performance?
@@
To enhance the noise robustness of BEV models, this work proposes a novel paradigm, NRSeg, a noise-resilient learning framework for BEV semantic segmentation, as shown in Fig. LABEL:Fig.intro_new(a). The proposed method simultaneously explores optimized guidance for synthetic data of varying quality and robust learning of the model itself against disturbances. A Perspective-Geometry Consistency Metric (PGCM) is designed to quantitatively assess the contribution of synthetic data to model learning. From the perspective view, the contribution scores of synthetic data are quantified by evaluating the similarity between road masks in synthetic images and their BEV back-projected counterparts. To alleviate the effect of generation noise in synthetic data, this score adaptively encourages the model to focus on non-labeled regions of synthetic data by guiding the loss optimization direction. Furthermore, the model’s robustness against disturbances is enhanced through dual-distribution constraints on predictions in the Bi-Distribution Parallel Prediction (BiDPP) module. While the multinomial distribution directly predicts semantic probabilities of a single pixel, the Dirichlet distribution is employed to model the uncertainty relationships based on Evidential Deep Learning (EDL) [30, 28]. The premise of EDL requires mutually exclusive semantic categories, which cannot be satisfied in BEV tasks. To effectively adapt EDL for BEV semantic segmentation tasks, the Hierarchical Local Semantic Exclusion (HLSE) module is designed, which partitions locally mutually exclusive semantic categories into unit clusters and achieves uncertainty modeling through hierarchical fusion of their independent learning processes.
@@ II-A BEV Semantic Segmentation
BEV semantic mapping performs semantic segmentation of static objects in road environments from a Bird’s Eye View (BEV). The fundamental challenge of BEV semantic segmentation lies in inferring 3D information from 2D inputs. Depending on the approach to view transformation, this research currently encompasses three distinct methodological categories. One dominant approach, exemplified by the LSS framework [29, 13], achieved spatial transformation through 2D pixel depth prediction. The alternative paradigm adaptively learned spatial feature representations by projecting 3D sampling points onto 2D image planes [21, 44]. Furthermore, some approaches [17, 15] leveraged the prior knowledge that static road objects predominantly lie on the ground plane, utilizing this prior information to facilitate view transformation through Inverse Perspective Mapping (IPM). While these view transformation methods demonstrate efficient adaptation to diverse road environments, their performance significantly degrades in complex, congested scenarios due to the sparsity of 2D observations [31]. To address this, StreamMapNet [49] introduced a streaming temporal strategy [37, 23] that compensated for single-frame sparsity by aggregating long-term historical information, improving mapping accuracy.
@@
Although these methods demonstrate competitive performance across domains, they have not yet explored directly enhancing the learning capability of the model by increasing the diversity of data distributions within the source domain. Benefiting from the rapid advancement of generative models [47, 51], it is now feasible to utilize simple yet effective information to synthesize diverse data. We aim to explore the potential of synthetic data from driving world models for robustifying BEV semantic segmentation.
@@ III-A Framework of the NRSeg Model
$Loss_{pv}=\text{SegLoss}(P_{v},L_{v}).\vskip-3.00003pt$ (1)
@@
$\vskip-3.00003ptF_{b}^{t^{\prime}}=\text{TFu}(\text{Ailgh}(F_{b}^{t-1}),F_{b}^{t}).$ (2)
@@
Moreover, $N$ is used to contain a shared BEV encoder and two segmentation detectors: a multinomial distribution $D_{m}^{k}$ and a Dirichlet distribution $D_{d}^{k}$ for predictions.
$\vskip-5.0ptD_{m}^{k},D_{d}^{k}=N(F_{b}^{t^{\prime}}),$ (5)
where $k$ is the number of semantic classes. The semantic probability predictions can be obtained by $P_{m}^{k}=\text{Sigmoid}(D_{m}^{k})$. Typically, the BEV semantic segmentation models are supervised by semantic labels $L_{b}$, which guide model optimization by constraining the predicted semantic probabilities $P_{m}^{k}$. Following the semi-supervised work [53], the DICE loss is adopted:
@@
In contemporary BEV perception datasets, perspective images and semantic labels are precisely aligned. Yet, due to generation errors in world models, synthetic data and semantic labels cannot maintain perfect correspondence. As shown in Fig. 1, to mitigate the misleading effects of such misalignment, this paper proposes the Perspective-Geometry Consistency Metric (PGCM) and Bi-Distribution Parallel Learning (BiDPP) to strengthen noise-resilient learning and model robustness in BEV semantic segmentation.
@@ III-B Perspective-Geometry Consistency Metric
First, the BEV semantic label and 3D object boxes are projected onto the perspective view, maintaining consistency with the process described in works [22, 19, 50]. Assuming the BEV label lies on the ground plane, the height is $h=0$. With this prior knowledge, each cell $(X,Y)$ of the BEV label $M_{b}$ can be projected into perspective views $(u,v)$:
$\vskip-3.00003pt[u,v,1]^{\text{T}}=\text{Proj}_{\text{b2p}}(T_{\text{ex}},T_{\text{in}},[X,Y,h]^{\text{T}}),$ (7)
where $T_{ex}$ and $T_{in}$ is the camera parameters. $Proj_{b2p}(\cdot)$ is the conversion from BEV coordinates to perspective coordinates. Then the projected road mask of perspective view $M_{o}$ can be obtained. The 3D object boxes are directly back-projected onto the perspective image through Eq. 7 to obtain the projected object mask $M_{c}$. The perspective view is a 2D representation, inherently suffering from occlusion issues due to viewpoint limitations. Since targets like vehicles exist above the ground plane, they occlude the road surface during perspective imaging. Consequently, the final reference perspective mask should adhere to the following criteria:
$\vskip-5.0ptM_{r}=M_{o}\cap\overline{M_{c}},$ (8)
where $\overline{M_{c}}$ is the negation of $M_{c}$. For synthetic images, perspective masks $M_{s}$ can be produced by pre-trained semantic segmentation models such as Mask2Former [4] and SAN [42].
To achieve accurate correspondence between synthetic data and BEV labels, the reference and synthetic perspective masks should inherently maintain geometric consistency. In other words, any misalignment between them indicates inherent discrepancies between the synthetic map and BEV semantic labels. Therefore, the alignment error between the reference perspective masks $M_{r}$ and the synthetic perspective masks $M_{s}$ is used to quantify the BEV accuracy of synthetic images. Considering that Intersection over Union (IoU) is a well-established metric for mask alignment evaluation, we adopt it as the primary measure here. The score of each frame oriented from synthetic data can be calculated by:
$\vskip-5.0pt\textbf{R}=\frac{1}{n}\sum_{i=0}^{n}\text{IoU}(M_{r}^{i},M_{s}^{i}),$ (9)
where $n$ is the number of perspective images in each frame. Then, for misaligned synthetic data samples and labels, this consistency score is incorporated as a weighting factor to guide the loss optimization direction in deep learning model training.
Synthetic data generated under BEV label control maintains overall structural consistency with the labels, with noise primarily manifesting as subtle structural shifts. Therefore, this paper selects the DICE loss [26], which favors holistic structural learning and exhibits high tolerance for minor errors. The DICE loss is defined as:
@@
where $H$ and $W$ are the shapes of labels. $P$ is the prediction probabilities and $L$ is the labels. Since the labels follow a binary distribution, Eq. 10 can be simplified as:
$\vskip-5.0ptloss_{gt}=1-\frac{2*P_{+}}{P_{+}+P_{-}+N_{+}+\sigma},$ (11)
@@
$\vskip-5.0pt\min_{p+,p_{-}}1-\frac{2*P_{+}}{P_{+}+P_{-}+N_{+}+\sigma}.$ (12)
Intuitively, this requires the numerator and denominator of the right-hand term in Eq. 12 to be equal. In other words, this enforces two conditions: (i) $P_{+}\approx N_{+}$, and (ii) $P_{-}\approx 0$.
@@
$\vskip-5.0ptloss_{gt}=1-\frac{2*P_{+}}{\textbf{R}*(P_{+}+P_{-}+N_{+}+\sigma)}.$ (13)
@@
$\vskip-5.0pt2*P_{+}\approx\textbf{R}*(P_{+}+P_{-}+N_{+}+\sigma).$ (14)
@@
$\vskip-5.0ptP_{-}\approx\frac{(2-\textbf{R})}{\textbf{R}}{\times}P_{+}-N_{+}-\sigma.$ (15)
@@ III-C Bi-Distribution Parallel Learning
In the BEV semantic segmentation task, the efficient multinomial distribution is typically adopted to predict semantic probabilities, where the loss function constrains the learning process. This constraint fundamentally relies on the reliability of the multinomial distribution assumption. In practice, however, the multinomial distribution exhibits inherent uncertainty when handling ambiguous categories, leading to potentially unreliable predictions. Specifically, the synthetic data referenced in this study inherently contains noise, which significantly increases the likelihood of unreliable predictions. Therefore, this section additionally introduces the Dirichlet distribution with Evidential Deep Learning (EDL) to further learn uncertainty. Consequently, it forms a parallel learning module together with the multinomial distribution, ensuring the robustness of the BEV semantic segmentation task.
Rooted in Dempster-Shafer Theory (DST) and Subjective Logic, EDL formalizes belief assignments through Dirichlet distributions, enabling principled uncertainty quantification [30]. Specifically, to learn a parametric representation of belief mass from the prediction, the Dirichlet distribution is employed as the prior for parameterization as $\alpha_{0},...,\alpha_{k}]$:
@@
where $e_{k}=\text{ReLU}(D_{d}^{k})$ is the evidence information, which is derived from the BEV semantic segmentation prediction $D_{d}^{k}$. Then, the parameterized distribution represents the density of probabilistic predictions $P_{e}$:
@@
where $S_{k}$ is the $k$-dimensional unit simplex. $B(\alpha)$ is a $k$-dimensional polynomial beta function. Theoretically, the expected probability for each BEV grid belonging to a certain semantic class corresponds to the mean of the Dirichlet distribution:
@@
where $S=\sum_{k}\alpha_{k}$. However, this theory requires that the semantic attributes of each pixel be mutually exclusive. Assuming two semantic attributes are non-exclusive, $\alpha_{1}\in(0,+\infty)$ and $\alpha_{2}\in(0,+\infty)$, their Dirichlet distribution means are:
@@
[…]
@@
Specifically, semantic classes are first hierarchically grouped based on their inherent properties, forming multi-level local clusters. For instance, in BEV semantic segmentation tasks, although drivable areas and lane markings are spatially non-mutually exclusive in terms of semantics, drivable areas and walkways are semantically mutually exclusive. Therefore, drivable areas and walkways can form a local cluster, $G_{m}$.
@@
Here, $n$ is the number of semantic classes in this cluster $G_{m}$ ($n<k$). Note that $n$ is not fixed to ensure adaptability across diverse tasks. Then, parameterization is implemented as the prior within each local cluster, $\alpha^{G_{m}}=[\alpha^{0},...,\alpha^{c_{n}}]$.
$\alpha^{G_{m}}=\text{RELU}(D_{d}^{k}[G_{m}])+1,$ (24)
where RELU is the active function. The evidence is $e^{G_{m}}=\text{RELU}(D_{d}^{k}[G_{m}])$. Similarly, predictive semantic probabilities and uncertainty measures are computed using the Dirichlet distribution. The difference is that uncertainty and probability are independently realized within each cluster $G_{m}$.
$P_{e}^{G_{m}}=\alpha^{G_{m}}/S^{G_{m}},\ \ \ S^{G_{m}}=\sum_{j=0}^{C_{n}}\alpha^{j}.$ (25)
To train this head, the type-II maximum likelihood version of the uncertainty-aware loss is adopted in the local clusters:
$Loss_{e}^{G_{m}}=\sum_{i=0}^{H_{b}{\times}W_{b}}\sum_{j=0}^{C_{n}}L_{i}^{j}\cdot log(S/\alpha_{i}^{j}).$ (26)
@@
[…]